OpenStudy (anonymous):

point 0,5 and 3,4 slope

OpenStudy (aroub):

4-5/ 3-0 -1/3

OpenStudy (saifoo.khan):

\[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

ooh nice subscripts

OpenStudy (saifoo.khan):

\[\frac{4-5}{3-0}\]\[\frac{-1}{3}\]

OpenStudy (saifoo.khan):

Thanks Sat.

OpenStudy (anonymous):

right 3, down one, slope is \[-\frac{1}{3}\]

OpenStudy (anonymous):

nice subscripts but mediocre fraction

OpenStudy (anonymous):

-\frac

OpenStudy (saifoo.khan):

Oh hmm.